OpenStudy (anonymous):

1. Solve 5x2 = –45x. using compleing the square

OpenStudy (mertsj):

First add 45x to both sides and then factor out the 5

OpenStudy (mertsj):

\[5(x^2+9x)=0\]

OpenStudy (mertsj):

Now complete the square by adding 81/4 inside the parentheses which is really adding 5 times 81/4 so now subtract 405/4 outside the parentheses

OpenStudy (mertsj):

\[5(x^2+9x+(\frac{9}{2})^2)-5(\frac{9}{2})^2=0\]

OpenStudy (mertsj):

\[5(x+\frac{9}{2})^2=5(\frac{9}{2})^2\]

OpenStudy (mertsj):

divide by 5 and then take the square root of both sides.

OpenStudy (mertsj):

Then subtract 9/2 from both sides. You will get 0 which is obviously the solution. Why you would be required to solve this equation by completing the square is beyond me.

OpenStudy (precal):

I agree I would have used an easier method.

OpenStudy (anonymous):

5x² = –45x 5x² + 45x = 0 5 (x² + 9x) = 0 5 ( x + 9/2)² = 5 * (9/2)² ->( x + 9/2)² = +/- (9/2) -> x + 9/2 = 9/2 => x = 0 , x = 9

OpenStudy (anonymous):

Oops, x = 0, x = -9